The Swifty Inc., a manufacturer of low-sugar, low-sodium, low-cholesterol TV dinners, would like to increase its market share in the Sunbelt. In order to do so, Swifty has decided to locate a new factory in the Panama City area. Swifty will either buy or lease a site depending upon which is more advantageous. The site location committee has narrowed down the available sites to the following three very similar buildings that will meet their needs. Building A: Purchase for a cash price of $612,100, useful life 26 years. Building B: Lease for 26 years with annual lease payments of $71,490 being made at the beginning of the year. Building C: Purchase for $655,200 cash. This building is larger than needed; however, the excess space can be sublet for 26 years at a net annual rental of $6,850. Rental payments will be received at the end of each year. The Swifty Inc. has no aversion to being a landlord. In which building would you recommend that The Swifty Inc. locate, assuming a 11% cost of funds? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

Net Present Value

Building A

$enter a dollar amount rounded to 0 decimal places

Building B

$enter a dollar amount rounded to 0 decimal places

Building C

$enter a dollar amount rounded to 0 decimal places
Answer #1

14.0799 L 1.658782 655200- 6850 x Low = 655200 - ( 6850x9.4881) = 655200 - 58143 = $597057 Thus, the minimum will be selected